OpenStudy (tumblewolf):

Need help with Trig limits Lim sin^2 theta / theta^2 Theta-0

OpenStudy (tkhunny):

Have you considered that this limit, without the squares, is 1?

OpenStudy (tumblewolf):

Yes so do I just square the 1's? It seemed wa y too easy

zepdrix (zepdrix):

As tk suggested, consider this limit identity,\[\large\rm \color{orangered}{\lim_{x\to0}\frac{\sin x}{x}=1}\] So what can you say about your problem?\[\large\rm \lim_{x\to0}\frac{(\sin x)^2}{x^2}\quad=\quad\lim_{x\to0}\left(\frac{\sin x}{x}\right)^2\quad=\quad \left(\large\rm \color{orangered}{\lim_{x\to0}\frac{\sin x}{x}}\right)^2\]

zepdrix (zepdrix):

Bahh I typed a little slow there >.< hehe

OpenStudy (tumblewolf):

Thank you though! Can you help me with a couple others?

OpenStudy (tumblewolf):

@zepdrix ?

OpenStudy (tkhunny):

Separate post, please.